### v1

@@ 1 Main
Finally, we discuss coherence times of representative NV defects close to dislocations, and we start by evaluating their axial () and transverse () components of the ground-state zero-field splitting (ZFS) tensor (Fig. S13). The highest (lowest) value is found to be 3839.0 (1796.6) MHz, deviating by up to 44% from that of the bulk NV (3230.8 MHz). The lowering of the defect symmetry leads to large nonzero values for the defects close to dislocations, with the highest predicted value reaching 813.3 MHz. Hence, the zero‐field degeneracy of the states is lifted to form non‐degenerate “clock” levels with no first‐order magnetic moment. As a result, NV centers at dislocations exhibit first‐order‐insensitive transitions in weak magnetic fields (), known as “clock transitions”. We computed the Hahn‐echo coherence time for magnetic fields up to mT [Figs. 5 (a), S14, and S15] and observe that decreases with increasing , reflecting enhanced decoherence from the field‐dependent dynamics of the 13C spin bath as one moves away from the zero‐field clock transition. This behavior is notably different from that of bulk NV (whose component of the ZFS is zero), which monotonically increases with and saturates at 1 ms [Fig 5 (a)] [32].
@@
The representative NV centers near dislocations exhibit longer coherence times than the bulk NV, at 0 mT up to 1 mT depending on the ZFS parameters (Figs. S14 and S15). The observed enhancement of at is determined by both the and components of the ZFS tensor. As illustrated in Fig. 5 (b), an increase in and leads to a significant enhancement in , reaching up to an order of magnitude improvement compared to bulk NV. A similar enhancement is also observed in the free induction decay time (Fig. S16). To further investigate the protection of spin coherence from environmental noise, we simulated NV coherence as a function of total evolution time under periodic Carr-Purcell-Meiboom-Gill (CPMG) pulse sequences at on representative NVs near dislocations. As shown in Fig. 5 (c), the coherence time of the - NV at the dislocation increases from 3.7 ms to 100 ms with -pulses, and the traces in Fig. 5 (d) show that raising suppresses decoherence and steepens the decay. We note that the reported values were computed in the long- limit, where longitudinal relaxation does not affect NV coherence [33, 34]. Our results demonstrate that the ZFS tensor of the NV can be engineered via dislocations, significantly enhancing coherence time and providing optimal working points for NV qubits.
Discussion
@@
We investigated the feasibility of using NV centers as spin qubits close to dislocations in diamond. We conducted a high-throughput screening of all possible NV configurations near the core regions of two types of dislocations, and showed that multiple configurations near the core can be stabilized in a triplet spin state and charge state. We then selected representative NVs and thoroughly examined their single-particle defect levels, optical cycle, and coherence times. In spite of differences in these properties relative to the bulk NVs, we found representative spin defects in the core region that meet the requisites to be useful qubits. Importantly, the optical cycle for qubit initialization and readout is feasible for several NVs. Moreover, certain centers can outperform the bulk NV, exhibiting a larger DWF and improved coherence time.
@@
We emphasize that the optical cycle of spin defects close to dislocations has never been investigated before; we could address this problem here thanks to the integration of a comprehensive suite of accurate and scalable simulation techniques that have only recently been developed [17, 18, 21, 23], enabling the investigation of the interplay between extended and point defects. Importantly, the workflow established here can be readily extended to explore other defects and hosts. We also note that given the large system size required to account for both point and extended defects (we used a 1728-atom supercell of diamond), our massively parallel GPU-accelerated implementation of TDDFT and QDET in the WEST code [21, 23, 35, 36] turned out to be critical to obtain accurate results, as was our recently developed workflow [17] to predict ISC rates.
@@
Based on the promising properties found in our study, we propose that NVs close to dislocations may be utilized to pattern and couple 1D arrays of spin defects; this engineering strategy to form quantum interconnects may advance the development of solid-state quantum technologies. Such a strategy is not limited to the system of NV qubits in diamond and could be utilized for other hosts such as SiC and other geometries, including semiconductor/oxide interfaces [37, 38].
@@
However, clearly some challenges remain to be addressed to fulfill the potential of the engineering strategy proposed here. Only a fraction of spin defects close to dislocations display the desired quantum properties. Hence, a controlled synthesis and manipulation of spin defects with favorable functionalities need to be developed. Part of the synthetic strategy may include the controlled introduction and patterning of specific dislocations into the sample, as well as the use of impurities (e.g., N) to engineer the desired Fermi level values.
Methods
@@
Using the 1728-atom supercells, we carried out a high-throughput search of configurations of NV centers close to dislocations using DFT as implemented in the Quantum ESPRESSO code [39]. We then computed their electronic and coherence properties, and compared them with those of NVs in bulk diamond. Using DFT, we computed the structural properties of spin defects with the PBE [26] functional, and formation energies in varied charge states, charge transition levels, and single-particle energies with the DDH functional [27]. We computed many-body energy levels using (spin-flip) TDDFT with the DDH functional, as implemented in the WEST code [21]. ISC rates were obtained with a newly developed framework [17] that combines calculations of many-body states with QDET [22, 23], fully relativistic calculations of SOC, and calculations of overlap functions using first-principles phonons [40, 41]. We used a seven-state model for state population simulations of ODMR spectra. Coherence times were obtained with the generalized cluster expansion techniques (gCCE) [24, 25] and the PyCCE code [42], and calculations of ZFS parameters were carried out with the PyZFS code [43]. Details are provided in the SI.
Supplementary Information Supplementary Text, Figs. S1–S16, Tables S1–S4 and References.
@@
Acknowledgements We thank F. Joseph Heremans, Alex High, Benchen Huang, and Marquis M. McMillan for useful discussions. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-05CH11231. This research also used resources of the Argonne Leadership Computing Facility at Argonne National Laboratory, a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-06CH11357. This work was supported by the AFOSR Grant No. FA9550-23-1-0330. Software development (WEST, PyCCE, and PyZFS) in this work was supported by MICCoM, as part of the Computational Materials Sciences Program funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences, and Engineering Division through Argonne National Laboratory.

### v2

@@ 2.4 Coherence times
Finally, we discuss coherence times of representative NV defects close to dislocations, and we start by evaluating their axial () and transverse () components of the ground-state zero-field splitting (ZFS) tensor (Fig. S13). The highest (lowest) value is found to be 3839.0 (1796.6) MHz, deviating by up to 44% from that of the bulk NV (3230.8 MHz). The lowering of the defect symmetry leads to large nonzero values for the defects close to dislocations, with the highest predicted value reaching 813.3 MHz. Hence, the zero‐field degeneracy of the states is lifted to form non‐degenerate “clock” levels with no first‐order magnetic moment. As a result, NV centers at dislocations exhibit first‐order‐insensitive transitions in weak magnetic fields (), known as “clock transitions”. We computed the Hahn‐echo coherence time for magnetic fields up to mT [Figs. 5 (a), S14, and S15] and observe that decreases with increasing , reflecting enhanced decoherence from the field‐dependent dynamics of the 13C spin bath as one moves away from the zero‐field clock transition. This behavior is notably different from that of bulk NV (whose component of the ZFS is zero), which monotonically increases with and saturates at 1 ms [Fig 5 (a)] [zhao_decoherence_2012].
@@
The representative NV centers near dislocations exhibit longer coherence times than the bulk NV, at 0 mT up to 1 mT depending on the ZFS parameters (Figs. S14 and S15). The observed enhancement of at is determined by both the and components of the ZFS tensor. As illustrated in Fig. 5 (b), an increase in and leads to a significant enhancement in , reaching up to an order of magnitude improvement compared to bulk NV. A similar enhancement is also observed in the free induction decay time (Fig. S16). To further investigate the protection of spin coherence from environmental noise, we simulated NV coherence as a function of total evolution time under periodic Carr-Purcell-Meiboom-Gill (CPMG) pulse sequences at on representative NVs near dislocations. As shown in Fig. 5 (c), the coherence time of the - NV at the dislocation increases from 3.7 ms to 100 ms with -pulses, and the traces in Fig. 5 (d) show that raising suppresses decoherence and steepens the decay. We note that the reported values were computed in the long- limit, where longitudinal relaxation does not affect NV coherence [balasubramanian_ultralong_2009, abobeih_one_2018]. Our results demonstrate that the ZFS tensor of the NV can be engineered via dislocations, significantly enhancing coherence time and providing optimal working points for NV qubits.
3 Discussion
@@
We investigated the feasibility of using NV centers as spin qubits close to dislocations in diamond. We conducted a high-throughput screening of all possible NV configurations near the core regions of two types of dislocations. Our formation energy calculations indicate that NVs are energetically attracted to dislocation cores, suggesting that dislocations can serve as a scaffold to position NVs in a 1D arrangement. We further showed that multiple NV configurations near the core can be stabilized in a triplet spin state and charge state. We then selected representative NVs and thoroughly examined their single-particle defect levels, optical cycle, and coherence times. In spite of differences in these properties relative to the bulk NVs, we found representative spin defects in the core region that meet the requisites to be useful qubits. Importantly, the optical cycle for qubit initialization and readout is feasible for several NVs. Moreover, certain centers can outperform the bulk NV, exhibiting a larger DWF and improved coherence time.
@@
We emphasize that the optical cycle of spin defects close to dislocations has never been investigated before; we could address this problem here thanks to the integration of a comprehensive suite of accurate and scalable simulation techniques that have only recently been developed [jin_first_2025, yu_gpu_2024, jin_excited_2023, chen_advances_2025], enabling the investigation of the interplay between extended and point defects. We also note that given the large system size required to account for both point and extended defects (we used a 1728-atom supercell of diamond), our massively parallel GPU-accelerated implementation of TDDFT and QDET in the WEST code [jin_excited_2023, chen_advances_2025, yu_gpu_2022, gavini_roadmap_2023] turned out to be critical to obtain accurate results, as was our recently developed workflow [jin_first_2025] to predict ISC rates.
@@
Based on the promising properties identified in our study, specifically the favorable energetics and the preserved or enhanced qubit properties, we propose that NVs close to dislocations can be utilized to pattern and couple 1D arrays of spin defects, forming the conceptual basis for dislocation-based quantum interconnects. This engineering strategy to create quantum interconnects may advance the development of solid-state quantum technologies. Moreover, the approach is not limited to NV qubits in diamond and could be applied to other hosts such as SiC and other geometries, including semiconductor/oxide interfaces [son_developing_2020, zhang_optical_2024].
@@
However, clearly, some challenges remain to be addressed to fulfill the potential of the engineering strategy proposed here. Only a fraction of spin defects close to dislocations display the desired quantum properties. Hence, a controlled synthesis and manipulation of spin defects with favorable functionalities needs to be developed. Part of the synthetic strategy may include the controlled introduction and patterning of specific dislocations into the sample, as well as the use of impurities (e.g., N) to engineer the desired Fermi level values.
Recently, experimental efforts have increasingly explored NV centers in low-symmetry environments as a means to tune and engineer their properties. Quasi-1D arrays of NVs have been created along heavy-ion tracks in diamond [liu_optical_2025], and NVs formed within diamond nanopillars have been demonstrated [kim_scalable_2025]. Additionally, NVs near surfaces [hughes_strongly_2025] and interfaces [lo_enhancement_2025] have been reported. The workflow established here can be readily extended to these systems, providing valuable insights to help interpret and guide experimental investigations in this regime.
4 Methods
4.1 Details of first-principles calculations
@@
Using the 1728-atom supercells, we carried out a high-throughput search of configurations of NV centers close to dislocations using DFT as implemented in the Quantum ESPRESSO code [giannozzi_quantum_2020]. We then computed their electronic and coherence properties, and compared them with those of NVs in bulk diamond. Using DFT, we computed the structural properties of spin defects with the PBE [perdew_generalized_1996] functional, and formation energies in varied charge states, charge transition levels, and single-particle energies with the DDH functional [skone_selfconsistent_2014]. We computed many-body energy levels using (spin-flip) TDDFT with the DDH functional, as implemented in the WEST code [jin_excited_2023]. ISC rates were obtained with a newly developed framework [jin_first_2025] that combines calculations of many-body states with QDET [sheng_green_2022, chen_advances_2025], fully relativistic calculations of SOC, and calculations of overlap functions using first-principles phonons [jin_photoluminescence_2021, jin_vibrationally_2022]. We used a seven-state model for state population simulations of ODMR spectra. Coherence times were obtained with the generalized cluster expansion techniques (gCCE) [yang_quantum_2008, onizhuk_probing_2021] and the PyCCE code [onizhuk_pycce_2021], and calculations of ZFS parameters were carried out with the PyZFS code [ma_pyzfs_2020]. Details are provided in the SI.
5 Data availability
Data that support the findings of this study are openly available at https://paperstack.uchicago.edu.
6 Code availability
Codes are openly available in public repositories.
Quantum ESPRESSO: https://gitlab.com/QEF/q-e
7 Acknowledgements
@@
We thank F. Joseph Heremans, Alex High, Benchen Huang, and Marquis M. McMillan for useful discussions. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-05CH11231. This research also used resources of the Argonne Leadership Computing Facility at Argonne National Laboratory, a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-06CH11357. This work was supported by the AFOSR Grant No. FA9550-23-1-0330. Software development (WEST, PyCCE, and PyZFS) in this work was supported by MICCoM, as part of the Computational Materials Sciences Program funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division through Argonne National Laboratory.
8 Author contributions
M.G. and G.G. conceived the research ideas. C.Z., V.W.-z.Y., Y.J., J.N., and S.P.G. performed the calculations. All authors contributed to the data analysis and manuscript writing.
9 Competing interests
The authors declare no competing financial or non-financial interests.
Supplementary Text, Figs. S1–S16, Tables S1–S4 and References.